Prepare two exhibits similar to Exhibit 11-6 and Exhibit 11-8, which show the following variances

Calgary Paper Company produces paper for photocopiers. The company has developed standard overhead rates based on a monthly capacity of 95,000 direct-labor hours as follows: 

Standard costs per unit (one box of paper): 

Variable overhead (2 direct-labor hours @ $6)     $12

Fixed overhead (2 direct-labor hours @ $6)           12

Total                                                              $24


During April, 40,000 units were scheduled for production: however, only 34,000 units were actually produced. The following data relate to April. 

1. Actual direct-labor cost incurred was $1,820,000 for 70,000 actual hours of work. 

2. Actual overhead incurred totaled $948,000, of which $448,000 was variable and $500,000 was fixed. 


Required: 

Prepare two exhibits similar to Exhibit 11-6 and Exhibit 11-8, which show the following variances. State whether each variance is favorable or unfavorable, where appropriate

 1. Variable-overhead spending variance.

 2. Variable-overhead efficiency variance.

 3. Fixed-overhead budget variance.

 4. Fixed-overhead volume variance. 
Answer #1

variable-overhead spending and efficiency variances

(hours = direct-labor hours)

(1)

(2)

(3)

(4)

Actual variable overhead

Projected variable overhead

Flexible budget: variable overhead

Variable overhead applied to work-in-process

Actual Qty (AQ)

x

Actual rate (AVR)

Actual Qty (AQ)

x

Standard rate (SVR)

Standard allowed Qty (SQ)

x

Standard rate (SVR)

Standard allowed Qty (SQ)

x

Standard rate (SVR)

70000

6.40

70000

6

68000

6

68000

6

Hours

Per hour

Hours

Per hour

Hours

Per hour

Hours

Per hour

448000

420000

408000

408000

$28000 U

$12000 U

No difference

Variable overhead spending variance

variable overhead efficiency variance

448000/7000 = 6.40

34000*2 = 68000

Fixed-overhead budget and volume variances

(hours = direct-labor hours)

(1)

(2)

(3)

Actual fixed overhead

Budgeted fixed overhead

Fixed overhead applied to work-in-process

Standard allowed Qty (SQ)

x

Standard fixed overhead rate

68000

6

Hours

Per hour

500000

480000

408000

$20000 U

$72000 U

Fixed overhead budget variance

Fixed overhead volume variance

40000*12 = 480000

34000*2 = 68000